OpenStudy (anonymous):

Use Eulers formula to find the missing number Vertices: 13 Edges: 28 Faces: ? a. 17 b. 16 c. 18 d. 20

OpenStudy (anonymous):

use V +F =E+2 here V=13 E=28 so F= 17

OpenStudy (anonymous):

Thank you so much! Would you mind helping me with a few more? :)

OpenStudy (anonymous):

yw